next up previous contents
Next: CDR Task Force: Up: Minutes of the Previous: Status of the

Status of Experiments

See hardcopies -- annex 7

GOLF: A. Gabriel

Off-set pointing (OSP) potential problem. Results of OSP available in 7-10 days.

VIRGO: C. Fröhlich

Need to exchange 292 faulty British Aerospace supplied capacitors.
Solar tests performed at Davos.
Science meeting to be held in March.

MDI: P. Scherrer

CCD contamination detected. Thomas: source is likely Silicon.
Gradient in continuum image detected, caused by polarizing beam splitter. Beam splitter will be exchanged.
Instrument has successfully measured P-modes from the ground.

SUMER: U. Schuehle

All requirements met with XDL detectors, except total count rate.
More power needed with XDL.
Problems with telescope and grating focus mechanism.

CDS: R. Harrison

Ground testing being performed.
Serious violation of 3" (arcsecond) resolution criterion for Zeiss mirror: 7" measured in spatial dimension.
With this resolution 40% scatter contribution expected in coronal hole, so valid science can still be done.

EIT: J-P Delaboudiniere

Problem with leaky door. Tightness to be verified in vibration test. Refurbishment of two qualification motors.

XDL Detectors: D. Machi

No schedule change for XDL since M-CDR review in Nov. '93. Delivery scheduled for Sept., Oct. '94.

UVCS: J. Kohl

Problem in OVI spectrometer: low efficiency in grating.
Roll rate reduction might reduce induced jitter to acceptable level. Study needed.
Spartan-201-1 data on polar coronal hole successful.

LASCO: D. Michels

Problems with 6 flight motors and doors.
Schedule getting very tight.

SWAN: J-L Bertaux

SWAN cannot meet ESA integration schedule. MMS delivery for integration expected by end of April '94.
Overall instrument efficiency factor 3 too low. Cause still unknown.

CELIAS/SEM: H. Hovestadt

STOF FM unit delivery delayed by more than one month, due to technological problems and illness of key personnel.

COSTEP: H. Kunow

Problems: (1) outgassing of p-tolualsulfonamide. Source needs to be identified and removed before S/C testing.
(2) EPHIN Si(Li) semiconductors need be replaced. FS not available until June '94.

ERNE: J. Torsti

ESU delivered without mechanical qualification due to failure in vibration test. Must be returned in Feb. for qualification tests.
CEPAC TV contamination problem: see COSTEP report.

next up previous contents
Next: CDR Task Force: Up: Minutes of the Previous: Status of the

Bernhard Fleck
Tue Jan 6 15:25:37 EST 1998